OpenStudy (anonymous):

If Dr. Suzuki rides his bike to his office, he averages 9 miles per hour. If he drives his car, he averages 36 miles per hour. His time driving 1/3 hour less than his time bicycling. How far is his office?

OpenStudy (anonymous):

Let the time for Dr Suzuki to ride his bike to his office be x 9x = 36 (x -1/3) 9x = 36x - 12 27x = 12 x = 12/27 9 * 12/27 = 4 miles

OpenStudy (karatechopper):

4 miles

hero (hero):

Yeah, just repeat what someone else says.

OpenStudy (karatechopper):

:(

hero (hero):

You would have been better off if you showed your steps

OpenStudy (karatechopper):

money did it for me

hero (hero):

No, you show your own steps, not his. He also posted an answer afterwards. That didn't stop you from posting one.

OpenStudy (karatechopper):

but it was the same steps! and thts not nice to say hero

OpenStudy (anonymous):

Hero, i see you getting rising quickly to the level of sensi, i dont know how your doing it but keep up the good work

hero (hero):

I'm surprised to hear you say that Lagrange. I thought you were going to say something mean.

OpenStudy (karatechopper):

y would he do tht?

OpenStudy (anonymous):

\[t \text { =} \frac{d}{r}\]Solve the following for d, the distance to the office:\[\frac{d }{9}=\frac{d}{36}+\frac{1}{3}\]d = 4 miles

OpenStudy (anonymous):

Thank you for the medal.

OpenStudy (karatechopper):

LOL